OpenStudy (anonymous):

Thanks for everyone that helped me I might have more questions later, but so far thanks

OpenStudy (anonymous):

np ^^

OpenStudy (anonymous):

what you need?

OpenStudy (anonymous):

Should I tell you guys now! I thought You might wanna break XD

OpenStudy (anonymous):

zoe has a bag of three differnt kinds of cookies 2/9 of the cookies have chocolate chips and 1/3 of the cookies have oatmeal and raisins how many of the cookies have chocolate chips or oatmeal and raisins? 1/9 2/27 5/9 2/3? I think its 5/9

OpenStudy (anonymous):

I don't stand the question,2/9 the total? and 1/2 is oatmeal + raisins ?

OpenStudy (anonymous):

no, 2/9ths are chocolate chip ;) and 1/3 of the cookies are oatmeal and raisins

OpenStudy (anonymous):

(2/9)+(1/3), Denominator is 9 than you have (1/3)*3=3/9 (2/9)+(3+9)=5/9

OpenStudy (anonymous):

yes, you have righ answer

OpenStudy (anonymous):

thanks